Question 26:

Ben Toucan, owner of the Aspen Restaurant, wants to determine the present value of his investment. The Aspen Restaurant is currently in the development stage but hopes to “begin” operations early next year. After-tax cash flows during the next five years are expected to be as follows: Year 1 = 0, Year 2 = 0, Year 3 = 0, Year 4 = $2.5 million, and Year 5 = $3 million. Cash inflows are expected to be $3.18 million in Year 6 and are expected to grow at a 6 percent annual rate thereafter. Recall that venture investors often use different discount rates when valuing ventures at various stages of their life cycles. For example, target discount rates by life cycle stage are development stage, 50 percent; startup stage, 40 percent; survival stage, 35 percent; and early rapid-growth stage, 30 percent. As ventures move from their late rapid-growth stages and into their maturity stages, a 20 percent discount rate is often used.

Questions:

- Determine the Aspen BrewPub’s terminal or horizon value at the end of five years.

VCF6 = VCF5 * (1+g) = ??

Terminal Value = VCF6 / (r-g) = ??

- What is the present value of the Aspen BrewPub?

PV = VCF1 / (1+r) + VCF2 / (1+r)^2 + ....+ VCF5 / (1+r)^5 + Terminal.value / (1+r)^5 = ??

- What percent ownership interest should Ben Toucan be willing to give to a venture investor, Sherri Isitar, for her $1,000,000 investment?

Ownership interest = ?
Answer #1

A). Used 20% discounting rate

VCF6 = VCF5 * (1+g) = $3 * (1.06)

= $3.18

Terminal Value = VCF6 / (r-g) = $3.18/(0.20 - 0.6)

= $ 22.71

B). Present Value Of Aspen Brew Pub

PV = VCF1 / (1+r) + VCF2 / (1+r)^2 + ....+ VCF5 / (1+r)^5 + Terminal.value / (1+r)^5 = ??

= $2.41+ $22.71/ 0.402

= $2.41 + $56.52

= $58.93

Year Cashflow PV factor @20% PV
1 $           -   0.833 $           -  
2 $           -   0.694 $           -  
3 $           -   0.579 $           -  
4 $       2.50 0.482 $       1.21
5 $       3.00 0.402 $       1.21
$       2.41

C. 58.93 % Stake